OpenStudy (anonymous):

Explain the difference between using the trigonometric ratios (sin, cos, tan) to solve for a missing angle in a right triangle versus using the reciprocal ratios (sec, csc, cot). You must use complete sentences and any evidence needed (such as an example) to prove your point of view. @poopsiedoodle @dan815

OpenStudy (dan815):

you would invert the ratios

OpenStudy (dan815):

sec=1/cos=1/(a/h)=h/a csc=1/sin=1/(o/h) = h/o

OpenStudy (dan815):